OpenStudy (anonymous):

how do you prove that lim as n-> infinity cos(cos...(cos(x))) exists where iterated cosine function is n times

OpenStudy (anonymous):

i.e. \[x_{1} = \cos (x) \] \[x_{n+1} = \cos(x_{n})\]